OpenStudy (anonymous):

An electric motor has an effective resistance of 32.0 ohms and an inductive reactance of 45.0 ohms when working underload. The rms voltage across the alternating source is 420 V. Calculate the rms current. (Answer: 7.61 A)

OpenStudy (anonymous):

R = 32 ohm XL = 45 ohm Z = sqrt(R^2 + XL^2) Z = sqrt(32^2 + 45^2) Z = sqrt(1024 + 2025) Z = sqrt(3049) Z = 55.21 I = V/Z = 420/55.21 I = 7.61 Ampere